OpenStudy (anonymous):

Please help, (2/(x+4))+1/(12/(x^2+9x+20))

OpenStudy (anonymous):

cross multiply ok

OpenStudy (anonymous):

ill set it up for you

OpenStudy (anonymous):



OpenStudy (goldphenoix):

You can use the equation button :| : \[\large \frac{ 2 }{ (x^2+9x+20 } = \frac{ 12 }{ (x+4) }\]

OpenStudy (anonymous):

2(x+4) distribute this and ill walk you through it

OpenStudy (anonymous):

@drehern u there